Determine the specific enthalpy of wet steam at a pressure of 70 kN m-2 and having a dryness fraction of 0.85.

1
Expert's answer
2021-04-30T11:22:51-0400

"H=C_p+\frac{pmr}{T}," "c_p=\frac{C_p}{m},"

"h=\frac Hm=c_p+\frac{pr}{T},"

"h=1007+\frac{70000\cdot 0.85}{393}=1158~\frac{J}{kg}."